MegaPrime Engineering
Status of MegaPrime after the December
2002 Engineering Run
-
Camera: working.
-
Shutter: working - rare intermittent fault on the axis card
-
Cryogenics: working.
-
Filter jukebox: kept inactive before undergoing upgrades in mid-January
-
Filters: only r' used
-
Prime Focus Environment: stable. Likely to have a few light leaks
as the jukebox is not tighten to the camera yet.
-
CUTE and cabling: functional.
-
Guiders/ISU: hardware working
-
Guiding software: Guide stars acquired on each guider, and guiding
done on one of the guiders with the ISU.
-
TCS: mount model and instrument zero done. Very limited communication
with MegaPrime Control already available.
-
Session software: working, including archiving.
-
Elixir and camera commissioning already done:
Zero point in r'
Noise check for the GFSU/FSA motors (OK!)
Mosaic geometry
Scattered light from nearby bright star (N/S done)
Darks on the telescope
Transfer curves on the telescope (no significant change with CEA data)
-
Focal plane / WFC alignment: One focal plane tilt already done -
will be further checked (though it seems to be a step in the right direction)